OpenStudy (anonymous):

Solve: x – 3y + z = 10 x + y – 3z = -10 -3x – y + z = 2 write your numeric answers as an ordered triple: (x,y,x)

OpenStudy (saifoo.khan):

Do you know how to start?

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

nope sorry i was helping my little brother. i have no clue how to do this at all

OpenStudy (saifoo.khan):

Okay. Read this carefully. Take 2 equations. Let's say 1 and 2. Solve them for "z". Equate them. (You'll get a polynomial without z then). Second, again take 2 equations. Let's say 2 and 3. Solve them for z. Equate them. (You'll get a polynomial without z then). Solve both of them simultaneously.

OpenStudy (anonymous):

ok can i show you my work to make sure i am doing it right?

OpenStudy (anonymous):

do i cancel the z's out so i multiply the equation 1 by 3 to get the z's the same to be able to cancel them.

OpenStudy (saifoo.khan):

No. That can get you wrong. Solve this way.

OpenStudy (saifoo.khan):

LOOK: x – 3y + z = 10 x + y – 3z = -10 -3x – y + z = 2 eq 1 and eq 2 z = 10 + 3y -x z= (x+y+10)/3 Do you get till here?

OpenStudy (anonymous):

kind of i understand you moved the z behind the = sign but how do you do that?

OpenStudy (saifoo.khan):

I did it in one step, mentally. I'll show you how i did that for equation 2 so you can get an idea: x + y – 3z = -10 y - 3z = -10 - x -3z = -10 -x -y 3z = 10 + x + y z = (10 + x + y)/3

OpenStudy (anonymous):

that makes since ok

OpenStudy (anonymous):

what do i do after that though

OpenStudy (saifoo.khan):

Okay. Now read my explanation again. I said: "Equate them. (You'll get a polynomial without z then)." z = 10 + 3y -x z= (x+y+10)/3 10 + 3y - x = (x+y+10)/3

OpenStudy (anonymous):

ok i got that so now what how do you zolve for it now?

OpenStudy (saifoo.khan):

Write it in the simplified form: 2y - 2x = -5 Now you can see that there's no z. NOW take 2 more equations and do all of this again.

OpenStudy (anonymous):

how did you simplify it like that though

OpenStudy (saifoo.khan):

10 + 3y - x = (x+y+10)/3 Multiply 3 on both sides and then isolate x and y on left side.

OpenStudy (anonymous):

that confuses me.....

OpenStudy (saifoo.khan):

Why? Where?

OpenStudy (anonymous):

like the whole simplifying step confuses me....... im really bad at this

OpenStudy (saifoo.khan):

Hold on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ooooo that makes since now thank you

OpenStudy (anonymous):

so can i show you my work to make sure i am doing it correctly

OpenStudy (anonymous):

i got 8x+2y=4 for the second one is that what you got?

OpenStudy (anonymous):

i dont know what to do after this or if this is even correct...... someone please help me